Hi Algebra Tutors, my name is Lauren and I just wanted to ask you a question . . . .
When there is an equation such as
50+x=89
My teacher always says that you have to find your answer and ALSO show all of your work as well showing all of the steps you did in the equation. So will you be able to help me with this? Email me back as soon as you get a chance, please.
Thanks A lot ,
Lauren =)

Answer by ptaylor(1894)

50+x=89
First, subtract 50 from each side. (equals subtracted from equals the differences are equal). This will isolate x by itself and that's what we want to do. When we isolate x by itself, we will then know what the value of x is.
50-50+x=89-50
On the left side, we know that 50-50=0 so now we only have x remaining on the left side.
0+x=39 When we subtract 50 from 89 on the right side, we have 39 left
So our answer is x=39
CK
substitute x=39 into the original equation and we have:
50+39=89
89=89

Hope this helps, Lauren---ptaylor